OpenStudy (eric_d):

A toy helicopter of mass 5.0kg rises with uniform acceleration from rest until a height of 60m in 10s. what is the upthrust produced by the helicopter rotor?

OpenStudy (eric_d):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

I'm not that great of a physics person. I would rather sit back and learn when someone walk this though with you. I think @Abhisar might know the answer and help you out.

OpenStudy (eric_d):

okay..

OpenStudy (eric_d):

@Abhisar

OpenStudy (abhisar):

Do u know the meaning of thrust ?

OpenStudy (eric_d):

nt sure

OpenStudy (eric_d):

upward force?

OpenStudy (abhisar):

It's basically nothing else than force !

OpenStudy (abhisar):

Yes upthrust means upward force which was provided inn a short duration of time.

OpenStudy (eric_d):

ok..

OpenStudy (mrnood):

@Abhisar It is simply the force - the time is not relevant in this case.

OpenStudy (abhisar):

Here, for helicopter u=0 s=60 t=10s

OpenStudy (mrnood):

The resultant force is upthrust - weight then acceleration = f/m you can work out the acceleration from the question, hence work out the resultant force, hence work out the thrust

OpenStudy (abhisar):

Can u calculate the acceleration ?

OpenStudy (eric_d):

v=u+at v=0+a(60)

OpenStudy (abhisar):